A sports car is advertised to have a maximum corne I need some help with this problem
Answer #1

a) We know, v^2 /r = centripetal acceleration a

V = sqrt(ar)

              = sqrt(0.85*9.8*100)

     = 28.86 m/s

b) v = sqrt(0.85*9.8*50)

          = 20.41 m/s

c)    (A)   v = sqrt(0.6*9.8*100) = 24.25 m/s

     (B) v = sqrt(0.6*9.8*50) = 17.15 m/s